(38) Number problems. You have at least $30 in change in your drawer, consisting of dimes and quarters. Write an inequality that shows the different number of coins in your drawer.
(54)Business and finance. If the inventor in exercise 53 charges $4 per unit, then her profit for producing and selling x units is given by the function.
P (x) = 2.25x –7000
(a) What is her profit if she sells 2000 units?
(b) What is her profit if she sells 5000 units?
(c) What is the break-even point for sales?

Answer by stanbon(48551)

You have at least $30 in change in your drawer, consisting of dimes and quarters. Write an inequality that shows the different number of coins in your drawer.
Let "d" be the number of dimes and "q" be the number of quarters.
Then the value of the dimes is 10d cents; Value of the quarters is 25q cents.
INEQUALITY:
10d+25q>=3000
-----------------
(54)Business and finance. If the inventor in exercise 53 charges $4 per unit, then her profit for producing and selling x units is given by the function.
P (x) = 2.25x –7000
(a) What is her profit if she sells 2000 units?
P(2000)= -$2500
(b) What is her profit if she sells 5000 units?
P(5000)=$4250
(c) What is the break-even point for sales?
The break-even point is P(x)=0
If 2.25x-7000=0
then x=7000/2.25=3111.11 units
Cheers,
Stan H.